Two point charges are on the y axis. A 6.00
Answer #1

The electric potential from each charge is K*q/r, where r is the distance to the charge measured from the origin. Potentials add algebraically. For the first charge, V1 = K*6.0*10^-6/0.0125 and for the second V2 = K*(-2.02*10^-6)/(-0.018). The total potential at the origin is V1 + V2.